### v1

@@ 2 Model
As is standard in learning theory, we assume a hypothesis class $\mathcal{H}$ of graphs, our goal is to design algorithms whose sample complexity is polynomial in the log size of the class and the inverses of the accuracy and confidence parameters. More specifically, we are given a (possibly huge) set $\mathcal{X}$ of nodes and a hypothesis class $\mathcal{H}$ of graphs on $\mathcal{X}$. We denote an unknown target graph $g^{\text{target}}$. The training set consists of a sequence $(x_{i},v_{i})_{i=1}^{m}$. The nodes $x_{1},\ldots,x_{m}$ are drawn IID from unknown distribution $\mathcal{D}$. For each node $x_{i}$, a random neighbor $v_{i}$ is drawn uniformly from its neighborhood $N_{g^{\text{target}}}(x_{i})$ in the target graph. The algorithm then outputs a graph $g^{\text{output}}$, and the goal is to minimize the expected precision and recall losses:
$\ell^{\textrm{precision}}(g^{\text{output}}):=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|N_{g^{\text{output}}}(x)\setminus N_{g^{\text{target}}}(x)|}{n_{g^{\text{output}}}(x)}\right]\,,$
$\ell^{\textrm{recall}}(g^{\text{output}}):=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|N_{g^{\text{target}}}(x)\setminus N_{g^{\text{output}}}(x)|}{n_{g^{\text{target}}}(x)}\right]\,,$
where for any graph $g$, $N_{g}(x)$ denotes the neighborhood of a node $x$ in $g$ and $n_{g}(x)=|N_{g}(x)|$ denotes the number of neighbors of $x$.
We focus on designing learning rules that can compete with the “best” graph in $\mathcal{H}$. Specifically, if there is a graph $g\in\mathcal{H}$ with precision and recall losses $p$ and $r$, respectively, can we output a graph $g^{\text{output}}$ whose precision and recall losses are comparable to $p$ and $r$? To answer this question we consider two natural metrics: scalar loss and Pareto loss.
@@
For any $\alpha>0$, we say $\alpha$-approximate optimal scalar loss is achievable if there exists a polynomial $P$ such that, for any finite hypothesis class $\mathcal{H}$, there is an algorithm $\mathcal{A}$ such that the following holds: For any $\varepsilon,\delta>0$ and any distribution $\mathcal{D}$, if $\mathcal{A}$ is given an IID training set of size $P(\log\lvert\mathcal{H}\rvert,1/\varepsilon,1/\delta)$, with probability at least $1-\delta$, it outputs a graph with scalar loss satisfying
@@
We emphasize that we aim to avoid dependencies on the number of vertices or edges in the graph, as allowing quadratic dependence on the number of nodes trivializes the problem. Our primary focus is on finite hypothesis classes, where we discuss dependencies on the cardinality of the hypothesis class, as is common in standard learning theory.
@@
Let $p,p^{\prime},r,r^{\prime}\in[0,1]$. We write $(p,r)\implies(p^{\prime},r^{\prime})$ to denote the following statement: there exists a polynomial $P$ such that, for any finite hypothesis class $\mathcal{H}$, there is an algorithm $\mathcal{A}$ such that the following holds: If $\mathcal{D}$ is a distribution for which there exists a graph in $\mathcal{H}$ with precision and recall losses $(p,r)$, then for any $\varepsilon,\delta>0$, if $\mathcal{A}$ is given $p,r$ and an IID training set of size $P(\log\lvert\mathcal{H}\rvert,1/\varepsilon,1/\delta)$, with probability at least $1-\delta$, it outputs a graph with precision and recall losses at most $p^{\prime}+\varepsilon$ and $r^{\prime}+\varepsilon$ [^4] Actually, our algorithms only requires the knowledge of $r$.
@@
This graph learning problem is considerably more challenging than standard supervised learning. If the entire neighborhood were observed in the training set rather than a random neighbor $v_{i}\sim\text{Unif}(N_{g^{\text{target}}}(x_{i}))$, the task would reduce to standard supervised learning. However, observing only a random neighbor prevents an unbiased estimate of precision loss, complicating the problem. We demonstrate it in the following example.
@@
In Fig 2, in the target graph, user $x_{i}$ likes a large number $n$ of songs. Our hypothesis class contains two graphs, each with exactly one edge connecting to user $x_{i}$. In the red graph graph $g_{1}$, the edge always connects to a true positive (i.e., the recommended song is one the user likes), while in the blue graph $g_{2}$, it connects to a false positive (i.e., the recommended song is one the user dislikes). Both graphs have high recall loss, $\ell^{\textrm{recall}}(g_{1})=\frac{n-1}{n}$ and $\ell^{\textrm{recall}}(g_{2})=1$. But graph $g_{1}$ has a precision loss of 00 as it always recommends a correct song, while graph $g_{2}$ has a precision loss of $1$ as it always recommends a wrong song. Since $n$ is large, it is unlikely that any song recommended by either graph will appear in the training set, making it impossible to distinguish between them, despite their drastically different precision losses.
$x_{i}$$u_{1}$$u_{2}$$u_{3}$$u_{n-1}$$u_{n}$$u^{\prime}$$\cdots$
Figure 2: The target graph (black) has neighborhood $N_{g^{\text{target}}}(x_{i})=\{u_{1},\ldots,u_{n}\}$ where $n$ is huge. The graph $g_{1}$ (red) has only one neighbor $u_{n}\in N_{g^{\text{target}}}(x_{i})$ while $g_{2}$ (blue) has only one neighbor $u^{\prime}\notin N_{g^{\text{target}}}(x_{i})$.
One might argue that the issue in the above example arises from the large degree of the target graph; however, we will later show that even when the target graph has a small degree, accurately estimating and optimizing precision remains impossible.
We emphasize that, unlike in other supervised learning settings, such as PAC learning, where minimizing empirical risk is often straightforward (e.g., by outputting any classifier that is consistent with the training set), here the learner only observes a single item $v_{i}$ per user $x_{i}$, rather than the user’s entire neighborhood. As a result, minimizing empirical precision loss in this context is far from trivial. For instance, regardless of the target graph, a complete graph is always consistent with the training set but can still incur high precision loss.
@@ 3 Main Results
We begin by presenting our results in the realizable setting, where the target graph belongs to the hypothesis class. In this case, there is no distinction between optimizing the scalar-loss objective and the Pareto-loss objective. We propose two new algorithms that achieve both the scalar-loss and Pareto-loss objectives.
@@
theoremrealizable In the realizable setting, there exist algorithms such that given an IID training set of size $m\geq O(\frac{\log(|\mathcal{H}|/\delta)}{\varepsilon})$, with probability at least $1-\delta$, the output graph $g^{\text{output}}$ satisfies
@@
Since $\mathds{1}(v_{i}\notin N_{g}(x_{i}))$ is an unbiased estimate of the recall loss $\ell^{\textrm{recall}}(g)$, any consistent graph (i.e., graph $g$ with $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin N_{g}(x_{i}))=0$) will have low recall loss. But ERM does not work as the training set contains only positive examples, and a complete graph is consistent with any training set but can incur high precision loss. Hence, the main challenge lies in minimizing the precision loss.
One of our proposed algorithms is based on the natural idea of maximum likelihood. At a high level, although multiple graphs may be consistent with the training set, for any observed edge $(x_{i},v_{i})$, if it is in graph $g$, the probability of observing this edge is $\frac{1}{n_{g}(x_{i})}$ when $g$ is the target graph. Consequently, we can rule out the complete graph, as its likelihood of generating any specific observed edge is low. The main challenge in the analysis, then, is how to connect the precision loss to the likelihood.
The other algorithm is more directly aligned with the scalar-loss objective. While we cannot obtain an unbiased estimate of the precision loss, and hence the scalar loss, we introduce a surrogate loss that both upper- and lower-bounds the scalar loss within a constant multiplicative factor. Then we output a graph minimizing this surrogate loss.
@@
In the agnostic setting, we show in the next two theorems that it is impossible to achieve an additive error for both scalar-loss and Pareto-loss objectives as is standard in learning theory. {restatable}theoremlowerscalar There exists a class =$\mathcal{H}=\{g_{1},g_{2}\}$ of two graphs, such that for any (possibly randomized improper) algorithm, there exists a target graph $g^{\text{target}}$ with bounded degree and a data distribution $\mathcal{D}$ with $\min_{g\in\mathcal{H}}(\ell^{\textrm{scalar}}(g))>0$ s.t. for any sample size $m>0$, with probability $1$ over the training set, the expected (over the randomness of the algorithm) loss of the output $g^{\text{output}}$
@@
theoremlowermulti There exists a class =$\mathcal{H}=\{g_{1},g_{2}\}$ of two graphs, such that for any (possibly randomized improper) algorithm given the knowledge of $(p,r)=(\frac{7}{16},\frac{1}{4})$, there exists a target graph $g^{\text{target}}$ with bounded degree and a data distribution $\mathcal{D}$ for which there exists a graph $g^{\dagger}\in\mathcal{H}$ with $\ell^{\textrm{recall}}(g^{\dagger})=\frac{1}{4}$ and $\ell^{\textrm{precision}}(g^{\dagger})=\frac{7}{16}$ s.t. for any sample size $m>0$, with probability $1$ over the training set, the expected (over the randomness of the algorithm) precision and recall losses of the output $g^{\text{output}}$ satisfy
@@
Hence the output graph either suffers $\ell^{\textrm{recall}}(g^{\text{output}})>\frac{1}{4}=\ell^{\textrm{recall}}(g^{\dagger})$ or $\ell^{\textrm{precision}}(g^{\text{output}})\geq\frac{23}{48}=\frac{23}{21}\ell^{\textrm{precision}}(g^{\dagger})$. Thus $(\frac{7}{16},\frac{1}{4})\not\Rightarrow(\frac{7}{16}+0.01,\frac{1}{4}+0.01)$.
Thus, in the scalar-loss case, we allow for a multiplicative factor $\alpha$. In the Pareto-loss case, we ask a more general question: which pairs of guarantees $(p^{\prime},r^{\prime})$ are achievable, given that there exists a hypothesis in the class with precision and recall losses $(p,r)$? Since the recall loss is optimizable, for any given $r$, if there exists a hypothesis in the class with recall loss $r$, we can always achieve that recall loss. Therefore, we refine our question as follows: given any $p,r\in[0,1]$, what is the minimum precision loss $p^{\prime}$ such that $(p,r)\Rightarrow(p^{\prime},r)$?
Since the recall is estimable, when we have an algorithm achieving $\alpha$-approximate scalar loss, we can first eliminate all graphs with recall loss higher than $r$ and then run this algorithm over the remaining graphs. Then we can achieve $(p,r)\Rightarrow(\alpha(p+r),r)$.
@@
theoremagnosticupper There exist an algorithm such that given an IID training set of size $m\geq O(\frac{\log(|\mathcal{H}|/\delta)}{\varepsilon^{2}})$, with probability at least $1-\delta$, the output graph $g^{\text{output}}$ satisfies
@@
This implies that for any $p,r\in[0,1]$, $(p,r)\Rightarrow(5(p+r),r)\,.$ This result is achieved using the same surrogate loss idea in the realizable setting. However, in the agnostic setting, applying maximum likelihood directly no longer works. This is because it is possible that none of the graphs in the hypothesis class are consistent with the training set and thus all graphs have zero likelihood. Instead, we make some modifications to adapt the maximum likelihood idea work for Pareto-loss objective.
@@
The results in the agnostic setting fail to offer meaningful guarantees in certain natural scenarios, such as $p=0$ and $r=\frac{1}{2}$ (i.e., when there exists a recommendation hypothesis that captures half of each user’s liked items without recommending any items the user does not like). We are therefore interested in the following question: whether $(p=0,r)\Rightarrow(p^{\prime}=0,r^{\prime}=r)$?
We propose an algorithm with sample complexity depending on the target graph’s degree and show that it is impossible to achieve zero precision loss with sample complexity independent of the target graph’s degree.
As discussed previously, it’s impossible for us to estimate the precision loss. However, we can still separate graphs with zero precision loss and non-zero precision loss if the target graph’s degree is bounded. If the precision loss
$\ell^{\textrm{precision}}(g,x)=\frac{|N_{g}(x)\setminus N_{g^{\text{target}}}(x)|}{n_{g}(x)}=1-\frac{|N_{g}(x)\cap N_{g^{\text{target}}}(x)|}{n_{g}(x)}$
of graph $g$ at user $x$ is 00, we have
$\frac{|N_{g}(x)\cap N_{g^{\text{target}}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}=\frac{1-\ell^{\textrm{precision}}(g,x)}{n_{g^{\text{target}}}(x)}=\frac{1}{n_{g^{\text{target}}}(x)}\,.$
@@
$\frac{|N_{g}(x)\cap N_{g^{\text{target}}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}=\frac{1-\ell^{\textrm{precision}}(g,x)}{n_{g^{\text{target}}}(x)}<\frac{1}{n_{g^{\text{target}}}(x)}\,.$
Hence, we can use the quantity $\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|N_{g}(x)\cap N_{g^{\text{target}}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}\right]$ to separate graphs with zero and non-zero precision loss, and it is estimable. For each graph $g$, when the gap of this quantity between graph with zero precision loss and $g$ is $\Delta_{g,\mathcal{D}}:=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{1}{n_{g^{\text{target}}}(x)}\right]-\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|N_{g}(x)\cap N_{g^{\text{target}}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}\right]>0$, then by obtaining $\frac{1}{\Delta_{g,\mathcal{D}}^{2}}$ samples of $x$, we can tell that $g$ has nonzero precision loss. Let $\Delta_{\mathcal{D}}$ be the smallest gap of this quantity between graphs with zero and nonzero precision losses:

### v2

@@ 2 Model
As is standard in learning theory, we assume a hypothesis class $\mathcal{H}$ of set functions, our goal is to design algorithms whose sample complexity is polynomial in the log size of the class and the inverses of the accuracy and confidence parameters. More specifically, we are given a input space $\mathcal{X}$, a label space $\mathcal{Y}$ and a hypothesis class $\mathcal{H}$, in which every hypothesis $h:\mathcal{X}\mapsto 2^{\mathcal{Y}}$ maps each $x$ to a set of labels in $\mathcal{Y}$. We denote an unknown target hypothesis $g^{\text{target}}$. The training set consists of a sequence $(x_{i},v_{i})_{i=1}^{m}$ with $(x_{i},v_{i})\in\mathcal{X}\times\mathcal{Y}$. The inputs $x_{1},\ldots,x_{m}$ are drawn IID from unknown distribution $\mathcal{D}$. For each input $x_{i}$, a random label $v_{i}$ is drawn uniformly from its ground-truth set, $g^{\text{target}}(x_{i})$ defined by the target hypothesis. The algorithm then outputs a hypothesis $g^{\text{output}}$, and the goal is to minimize the expected precision and recall losses:
$\ell^{\textrm{precision}}(g^{\text{output}}):=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|g^{\text{output}}(x)\setminus g^{\text{target}}(x)|}{n_{g^{\text{output}}}(x)}\right]\,$,
$\ell^{\textrm{recall}}(g^{\text{output}}):=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|g^{\text{target}}(x)\setminus g^{\text{output}}(x)|}{n_{g^{\text{target}}}(x)}\right]\,$,
where for any hypothesis $g$, $n_{g}(x)=|g(x)|$ denotes the set size of $g$ at $x$.
We focus on designing learning rules that can compete with the “best” hypothesis in $\mathcal{H}$. Specifically, if there is a hypothesis $g\in\mathcal{H}$ with precision and recall losses $p$ and $r$, respectively, can we output a hypothesis $g^{\text{output}}$ whose precision and recall losses are comparable to $p$ and $r$? To answer this question we consider two natural metrics: scalar loss and Pareto loss.
@@
For any $\alpha>0$, we say $\alpha$-approximate optimal scalar loss is achievable if there exists a polynomial $P$ such that, for any finite hypothesis class $\mathcal{H}$, there is an algorithm $\mathcal{A}$ such that the following holds: For any $\varepsilon,\delta>0$ and any distribution $\mathcal{D}$, if $\mathcal{A}$ is given an IID training set of size $P(\log\lvert\mathcal{H}\rvert,1/\varepsilon,1/\delta)$, with probability at least $1-\delta$, it outputs a hypothesis with scalar loss satisfying
@@
Our primary focus is on finite hypothesis classes, where we discuss dependencies on the cardinality of the hypothesis class, as is common in standard learning theory. We emphasize that the average loss is a natural way to combine precision and recall losses. Other scalarizations, such as the F1 score, are also possible. Several results in this paper extend to alternative definitions of scalar loss functions, which we discuss in the Discussion section.
@@
Let $p,p^{\prime},r,r^{\prime}\in[0,1]$. We write $(p,r)\implies(p^{\prime},r^{\prime})$ to denote the following statement: there exists a polynomial $P$ such that, for any finite hypothesis class $\mathcal{H}$, there is an algorithm $\mathcal{A}$ such that the following holds: If $\mathcal{D}$ is a distribution for which there exists a hypothesis in $\mathcal{H}$ with precision and recall losses $(p,r)$, then for any $\varepsilon,\delta>0$, if $\mathcal{A}$ is given $p,r$ and an IID training set of size $P(\log\lvert\mathcal{H}\rvert,1/\varepsilon,1/\delta)$, with probability at least $1-\delta$, it outputs a hypothesis with precision and recall losses at most $p^{\prime}+\varepsilon$ and $r^{\prime}+\varepsilon$. [^4] Actually, our algorithms only requires the knowledge of $r$.
@@
This hypothesis learning problem is considerably more challenging than standard supervised learning. If the entire ground-truth set $g^{\text{target}}(x_{i})$ were observed in the training set rather than a random label $v_{i}\sim\text{Unif}(g^{\text{target}}(x_{i}))$, the task would reduce to standard supervised learning. However, observing only a random label prevents an unbiased estimate of precision loss, complicating the problem. We demonstrate it in the following example.
@@
In Fig 2, the target hypothesis associates input $x_{i}$ with a large number $n$ of labels. Our hypothesis class contains two hypotheses, each predicting exactly one label for input $x_{i}$. In the red hypothesis $g_{1}$, the predicted label is always a true positive (i.e., it belongs to the ground-truth set), while in the blue hypothesis $g_{2}$, the predicted label is a false positive (i.e., it is not in the ground-truth set). Both hypotheses incur high recall loss: $\ell^{\textrm{recall}}(g_{1})=\frac{n-1}{n}$ and $\ell^{\textrm{recall}}(g_{2})=1$. However, they differ significantly in precision loss: $g_{1}$ has a precision loss of 0, as it always predicts a correct label, whereas $g_{2}$ has a precision loss of 1, as it always predicts an incorrect label. Since $n$ is large, the probability of observing either predicted label in the training data is very low, making it impossible to distinguish between these two hypotheses—despite their drastically different precision losses.
$x_{i}$$u_{1}$$u_{2}$$u_{3}$$u_{n-1}$$_{n}uu^{\prime}\cdots$
Figure 2: The target hypothesis (black) outputs $g^{\text{target}}(x_{i})=\{u_{1},\ldots,u_{n}\}$ where $n$ is huge. The hypothesis $g_{1}$ (red) outputs only one label $u_{n}\in g^{\text{target}}(x_{i})$ while $g_{2}$ (blue) outputs only one label $u^{\prime}\notin g^{\text{target}}(x_{i})$.
One might argue that the issue in the above example arises from the large size of the ground-truth set; however, we will later show that even when the ground-truth set has a small size, accurately estimating and optimizing precision remains impossible.
We emphasize that, unlike in other supervised learning settings, such as PAC learning, where minimizing empirical risk is often straightforward (e.g., by outputting any classifier that is consistent with the training set), here the learner only observes a single label $v_{i}$ per input $x_{i}$, rather than its entire ground-truth set. As a result, minimizing empirical precision loss in this context is far from trivial. For instance, regardless of the ground-truth set, a complete set function, which outputs the entire label space $\mathcal{Y}$ for any input, is always consistent with the training data but can still incur high precision loss.
@@ 3 Main Results
We begin by presenting our results in the realizable setting, where the target hypothesis belongs to the hypothesis class $\mathcal{H}$. In this case, there is no distinction between optimizing the scalar-loss objective and the Pareto-loss objective. We propose two new algorithms that achieve both the scalar-loss and Pareto-loss objectives: one based on maximum likelihood estimation, and the other on minimizing a surrogate loss.
@@
theoremrealizable In the realizable setting, there exist algorithms such that given an IID training set of size $m\geq O(\frac{\log(|\mathcal{H}|/\delta)}{\varepsilon})$, with probability at least $1-\delta$, the output hypothesis $g^{\text{output}}$ satisfies
@@
Since $\mathds{1}(v_{i}\notin g(x_{i}))$ is an unbiased estimate of the recall loss $\ell^{\textrm{recall}}(g)$, any consistent hypothesis (i.e., hypothesis $g$ with $\sum_{i=1}^{m}\mathds{1}(v_{i}\notin g(x_{i}))=0$) will have low recall loss. But ERM does not work as the training set contains only positive examples, and a complete set function is consistent with any training set but can incur high precision loss. Hence, the main challenge lies in minimizing the precision loss. Below are high-level descriptions of two algorithms designed to tackle this problem.
Algorithm 1: Maximum Likelihood.
One of our proposed algorithms is based on the natural idea of maximum likelihood. At a high level, although multiple hypotheses may be consistent with the training set, for any observed input-label pair $(x_{i},v_{i})$, if $v_{i}$ is contained in $g(x_{i})$, the probability of observing this label is $\frac{1}{n_{g}(x_{i})}$ when $g$ is the target hypothesis. The maximum likelihood method returns $g^{\text{output}}=\operatorname*{arg\,max}_{g\in\mathcal{H}}\prod_{i=1}^{m}\frac{\mathds{1}(v_{i}\in g(x_{i}))}{n_{g}(x_{i})}$, which is equivalent to returning the hypothesis with the minimum sum of log output sizes among all consistent hypotheses, i.e.,
Consequently, we can rule out the complete set function, as its sum of log output size $\sum_{i=1}^{m}\log(n_{g}(x_{i}))$ is huge. Any consistent hypothesis will have low empirical recall loss by applying standard concentration inequality and thus, $\widehat{\ell}^{\textrm{recall}}(g^{\text{output}})$ is small and we only need to show that the empirical precision loss is small. The main technical challenge in the analysis is how to connect precision loss with likelihood.
Algorithm 2: Minimizing a Surrogate Loss.
The other algorithm is more directly aligned with the scalar-loss objective. While we cannot obtain an unbiased estimate of the precision loss, and hence the scalar loss, we introduce a surrogate loss that both upper- and lower-bounds the scalar loss within a constant multiplicative factor. Then we output a hypothesis minimizing this surrogate loss.
For any hypothesis $g$, we define a vector $v_{g}:\mathcal{H}\times\mathcal{H}\to[0,1]$ by
$v_{g}(g^{\prime},g^{\prime\prime})=\frac{1}{m}\sum_{i=1}^{m}\mathbb{P}_{v\sim\mathrm{Unif}(g(x_{i}))}\left(v\in g^{\prime}(x_{i})\setminus g^{\prime\prime}(x_{i})\right).$
Intuitively, $v_{g}(g^{\prime},g^{\prime\prime})$ represents the fraction of correct labels that are output by $g^{\prime}$ but not by $g^{\prime\prime}$ in the counterfactual scenario where $g$ is the target hypothesis. If $g$ is indeed the target hypothesis, then this quantity should be consistent with our training data, $v_{g}(g^{\prime},g^{\prime\prime})\approx v_{\widehat{g}}(g^{\prime},g^{\prime\prime})$, where $\widehat{g}$ is the observed (empirical) hypothesis; i.e., the hypothesis in which every $x_{i}$ is only associated with the random label $v_{i}$ which is observed in the training set.
We then define a metric $d_{\mathcal{H}}$ between two hypotheses $g_{1}$ and $g_{2}$ by
$d_{\mathcal{H}}({g_{1},g_{2}})=\|v_{g_{1}}-v_{g_{2}}\|_{\infty}\,.$
Surprisingly, we show that $d_{\mathcal{H}}(g^{\text{target}},g)$ is a surrogate for the scalar loss, providing both lower and upper bounds on the scalar loss with a constant multiplicative factor.
@@
In the agnostic setting, we show in the next two theorems that it is impossible to achieve an additive error for both scalar-loss and Pareto-loss objectives as is standard in learning theory. {restatable}theoremlowerscalar There exists a class $\mathcal{H}=\{g_{1},g_{2}\}$ of two hypotheses, such that for any (possibly randomized improper) algorithm, there exists a target hypothesis $g^{\text{target}}$ with bounded output set size (that is, there exists a constant $C$ such that $|g^{\text{target}}(x)|\leq C$ for all $x\in\mathcal{X}$) and a data distribution $\mathcal{D}$ with $\min_{g\in\mathcal{H}}(\ell^{\textrm{scalar}}(g))>0$ s.t. for any sample size $m>0$, with probability $1$ over the training set, the expected (over the randomness of the algorithm) loss of the output $g^{\text{output}}$
@@
theoremlowermulti There exists a class $\mathcal{H}=\{g_{1},g_{2}\}$ of two hypotheses, such that for any (possibly randomized improper) algorithm given the knowledge of $(p,r)=(\frac{7}{16},\frac{1}{4})$, there exists a target hypothesis $g^{\text{target}}$ with bounded output set size and a data distribution $\mathcal{D}$ for which there exists a hypothesis $g^{\dagger}\in\mathcal{H}$ with $\ell^{\textrm{recall}}(g^{\dagger})=\frac{1}{4}$ and $\ell^{\textrm{precision}}(g^{\dagger})=\frac{7}{16}$ s.t. for any sample size $m>0$, with probability $1$ over the training set, the expected (over the randomness of the algorithm) precision and recall losses of the output $g^{\text{output}}$ satisfy
@@
Hence the output hypothesis either suffers $\ell^{\textrm{recall}}(g^{\text{output}})>\frac{1}{4}=\ell^{\textrm{recall}}(g^{\dagger})$ or $\ell^{\textrm{precision}}(g^{\text{output}})\geq\frac{23}{48}=\frac{23}{21}\ell^{\textrm{precision}}(g^{\dagger})$. Thus $(\frac{7}{16},\frac{1}{4})\not\Rightarrow(\frac{7}{16}+0.01,\frac{1}{4}+0.01)$.
For any hypothesis $g$, it’s precision loss at any input $x$ is $\frac{|g(x)\setminus g^{\text{target}}(x)|}{n_{g}(x)}$ and we only get a random label $v\sim\mathrm{Unif}(g^{\text{target}}(x))$. If we are given the knowledge of the output size $n_{g^{\text{target}}}(x)$ of the target hypothesis, then we can obtain an unbiased estimate of the precision loss, i.e., $1-\frac{n_{g^{\text{target}}}(x)}{n_{g}(x)}\cdot\mathds{1}(v\in g(x))$. But the difficulty lies in that we don’t know $n_{g^{\text{target}}}(x)$.
Based on these two lower bounds, in the scalar-loss case, we allow for a multiplicative factor $\alpha$. In the Pareto-loss case, we ask a more general question: which pairs of guarantees $(p^{\prime},r^{\prime})$ are achievable, given that there exists a hypothesis in the class with precision and recall losses $(p,r)$? Since the recall loss is optimizable, for any given $r$, if there exists a hypothesis in the class with recall loss $r$, we can always achieve that recall loss. Therefore, we refine our question as follows: given any $p,r\in[0,1]$, what is the minimum precision loss $p^{\prime}$ such that $(p,r)\Rightarrow(p^{\prime},r)$?
Since the recall is estimable, when we have an algorithm achieving $\alpha$-approximate scalar loss, we can first eliminate all hypotheses with recall loss higher than $r$ and then run this algorithm over the remaining hypotheses. Then we can achieve $(p,r)\Rightarrow(\alpha(p+r),r)$.
@@
theoremagnosticupper There exist an algorithm such that given an IID training set of size $m\geq O(\frac{\log(|\mathcal{H}|/\delta)}{\varepsilon^{2}})$, with probability at least $1-\delta$, the output hypothesis $g^{\text{output}}$ satisfies
@@
This implies that for any $p,r\in[0,1]$, $(p,r)\Rightarrow(5(p+r),r)\,.$ This result is achieved using the same surrogate loss idea in the realizable setting. However, in the agnostic setting, applying maximum likelihood directly no longer works. This is because it is possible that none of the hypotheses in the hypothesis class are consistent with the training set and thus all hypotheses have zero likelihood. Instead, we make some modifications to adapt the maximum likelihood idea work for Pareto-loss objective.
As mentioned earlier, the maximum likelihood method is equivalent to returning the hypothesis with the minimum sum of log output sizes among all consistent hypotheses. Hence, it can be decomposed into two steps: minimizing the recall loss and then regulating by minimizing the sum of log output sizes. In the agnostic setting, given any $r$, we first find the set $\widehat{\mathcal{H}}$ of all hypotheses in the hypothesis class with recall loss at most $r+2\varepsilon$, and then regulate by minimizing the sum of log truncated output sizes by returning
where $a\wedge b:=\min(a,b)$. The truncation plays an important role here. Intuitively, let $g^{\dagger}$ denote the hypothesis with precision and recall losses $p$ and $r$, respectively. If there exists an $x_{i}$ such that $n_{g^{\dagger}}(x_{i})$ is very large, minimizing the sum of log-untruncated output sizes will never return $g^{\dagger}$. By applying truncation, we limit the effect of a single input with a very large output size.
@@
The results in the agnostic setting fail to offer meaningful guarantees in certain natural scenarios, such as $p=0$ and $r=\frac{1}{2}$ (i.e., when there exists a hypothesis that captures half of ground-truth labels without outputting any incorrect labels). We are therefore interested in the following question: whether $(p=0,r)\Rightarrow(p^{\prime}=0,r^{\prime}=r)$?
We propose an algorithm with sample complexity depending on output set size of the target hypothesis and show that it is impossible to achieve zero precision loss with sample complexity independent of the target hypothesis’s output set size.
As discussed previously, it’s impossible for us to estimate the precision loss. However, we can still separate hypotheses with zero precision loss and non-zero precision loss if the the target hypothesis’s output set size is bounded. If the precision loss of hypothesis $g$ at input $x$ is 0, i.e.,
$\ell^{\textrm{precision}}(g,x)=\frac{|g(x)\setminus g^{\text{target}}(x)|}{n_{g}(x)}=1-\frac{|g(x)\cap g^{\text{target}}(x)|}{n_{g}(x)}=0\,$,
we have
$\frac{|g(x)\cap g^{\text{target}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}=\frac{1-\ell^{\textrm{precision}}(g,x)}{n_{g^{\text{target}}}(x)}=\frac{1}{n_{g^{\text{target}}}(x)}\,.$
@@
$\frac{|g(x)\cap g^{\text{target}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}=\frac{1-\ell^{\textrm{precision}}(g,x)}{n_{g^{\text{target}}}(x)}<\frac{1}{n_{g^{\text{target}}}(x)}\,.$
Hence, we can use the quantity $\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|g(x)\cap g^{\text{target}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}\right]$ to separate hypotheses with zero and non-zero precision loss, and it is estimable. For each hypothesis $g$, when the gap of this quantity between hypothesis with zero precision loss and $g$ is $\Delta_{g,\mathcal{D}}:=\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{1}{n_{g^{\text{target}}}(x)}\right]-\mathbb{E}_{x\sim\mathcal{D}}\left[\frac{|g(x)\cap g^{\text{target}}(x)|}{n_{g}(x)\cdot n_{g^{\text{target}}}(x)}\right]>0$, then by obtaining $\frac{1}{\Delta_{g,\mathcal{D}}^{2}}$ samples of $x$, we can tell that $g$ has nonzero precision loss. Let $\Delta_{\mathcal{D}}$ be the smallest gap of this quantity between hypotheses with zero and nonzero precision losses: